Solve each formula for the indicated letter. Assume that all variables represent non-negative numbers.

A = A0(1 – r)2 for r
(A business formula)

Answer by stanbon(75887)

A = A0(1 – r)2 for r
-------------------
This is probably what you meant:
A = Ao(1-r)^2
----
(1-r)^2 = A/Ao
----
1-r = +-sqrt[A/Ao]
----
r = 1+-sqrt[A/Ao]
=======================
Cheers,
Stan H.
========